Could someone give some guidance on how to solve c and d?1. During the last week of classes before finals, Ryan Rabbitt starts to find coffee a daily necessity Every morning, Ryan ma
Answer #1

As you mention only about the 3rd and 4th one, i'll go briefly for the first and 2nd one.

We can write  \small \frac{dV}{dt}=k(2 \pi r h+\pi r^2 h), where \small k is the proportionality constant. And dV .

So we get the differential equation as

dh 2k .

After solving this we get  h(t) as,

\small h(t)=\frac{r}{2k}(C e^{\frac{2k}{r}t}-k), where r is the radius of the cylinder and C is the unknown arbitrary number, yet to find out.

Now we have to unknown in the equation k and C, to find these we need to two equation. This conditions are given the 3rd problem. Initially the filter was full i.e ate t=0, h=12 cm and when t=2.2 min, h=0 as filtration process was over.

Now inputting the 1st condition in the previous equation we get

\small \frac{C}{k}=\frac{29}{5},

and putting the 2nd condition we get,

20, rearranging this equation and putting the value of \small \frac{C}{k} we get \small k=-\frac{5}{2} cm/min.

So \small C=-\frac{29}{2} cm/min.

Getting k a -ve number was quite normal because \small \frac{dV}{dt} becomes a -ve number which means as time increase volume will decrease.

For thew fourth part, as rate of volume is constant the filter tube will maintain a height of \small h=\frac{6}{\pi r^2}=\frac{6}{25 \pi} cm .